ahtzymackayla:

How would I start to solve 44=32+y?

Gdeinward:

You must isolate y.

Gdeinward:

to do that, you must eliminate 32 from both sides through subtraction

ahtzymackayla:

so I subtract 32 from 44?

Gdeinward:

44=32 + y 44-32= 32-32 + y 12 = y

Gdeinward:

correct

ahtzymackayla:

Oh okay

ahtzymackayla:

There's more problem similar to that one but now I know hot to solve the problems thank you :)

Gdeinward:

Of course, my pleasure.